# Quantum computing with atomic qubit arrays: confronting the cost of connectivity

### v1

@@ IV.2 Transport based gates
An alternative approach which enables arbitrary longer range connectivity is qubit motion during circuit operation. The concept of relying on physical motion of qubits dates back to early ideas for scaling trapped ion quantum processors[172, 173] and has been implemented with great success in recent trapped ion demonstrations[174, 175]. In trapped ion platforms, and also extensions to shuttling of quantum dots in semiconductor structures[176, 177], motion relies on applying a sequence of voltages to closely spaced electrodes to provide a time dependent force on the charged particle. A drawback of the trapped ion approach, in addition to the complexity of controlling many electrodes, is that motion induces substantial heating of the center of mass motional state so recooling is necessary as part of the control sequence.
@@
Neutral atoms trapped in optical potentials defined by light can be transported much more efficiently in terms of the number of required control signals: either by varying the phase of a single optical beam to continuously move the nodes of an interference pattern[178], or by scanning an optical tweezer using a varying frequency applied to an acousto-optical deflector[179]. When applied to moving arrays of atoms in parallel the optical tweezer approach becomes a powerful technique for connecting logical qubits in a neutral atom processor[19, 45, 47].
@@
We may compare the time to implement a logical CZ gate using atom motion with the case of long-range Rydberg interactions discussed in Sec. IV.1. The attainable transport speed is limited by the need to minimize motional heating[180, 181]. Several different transport profiles have been demonstrated including constant jerk[19], and minimal jerk[182, 183, 49], where jerk is the time derivative of acceleration[184]. Remarkably it is also possible to throw and catch atoms over distances of several $\mu\rm m$[185].
@@
For the minimal jerk profile we find using the analysis in [186] that the time to move a distance $R$ while incurring $\delta n$ quanta of motional excitation is[40]
@@
Figure 8: Sequence for a transport based logical $\sf CZ$ involves (top to bottom) moving the data qubits a distance $R$, applying a Rydberg pulse, and then moving the data qubits back. The contour plot shows the minimal jerk time to implement the move back and forth, neglecting any additional time for the Rydberg pulse, as a function of the distance $R=dr$, and the round trip motional heating incurred. The inset shows the one-dimensional atomic temperature as a function of the mean number of motional quanta assuming a trap frequency of $\omega_{0}=2\pi\times 100~{}\rm kHz$ and the expression $k_{\rm B}T=\hbar\omega_{0}/\ln(1+1/\langle n\rangle)$.
@@
Parallel implementation of a transport based transversal gate relies on positioning the atom pairs at a small spacing, relative to the distance to neighboring pairs. Doing so allows multiple gate pairs to be performed in the same time step. As long as the array pitch $r$, which is the distance between pairs, satisfies $r\gg r_{g}$, with $r_{g}$ the distance between atoms in one gate pair there will be minimal crosstalk. The sensitivity of the time-optimal $\sf CZ$ gate to interaction crosstalk was studied in [125] where it was found that provided $\eta_{i}=\sum_{j}\Delta_{ij}t_{\rm gate}/(2\pi)<0.01$ the gate infidelity due to crosstalk is upper bounded by $10^{-4}$. Here $\Delta_{ij}$ is the energy shift of Rydberg atom $i$ due to all other atoms $j$, and $t_{\sf CZ}$ is the gate time. In the van der Waals interaction limit the shifts scale as $\Delta_{ij}\simeq\frac{V(r_{g})}{\hbar}\left(\frac{r_{g}}{r_{ij}}\right)^{6}.$ For a high fidelity time-optimal gate we require[122] $\Omega t_{\sf CZ}/\hbar\geq 7.6$ with $\Omega$ the Rydberg excitation Rabi frequency. Using $(V/\hbar)/\Omega=20$ which is adequate for a high fidelity gate and approximating $\eta_{i}$ by the worst case of $\eta\simeq 4\Delta(r)t_{\sf CZ}/(2\pi)$ we find
@@
Using $r_{g}=1.4~{}\mu\rm m$ (the same tight array spacing as was assumed in Fig. 6) we require $r=6.44~{}\mu\rm m$ so the transport distance for a $d=10$ code is $R=64~{}\mu\rm m$. Referring to Fig. 8 the total transport time at a heating increment of $\delta n=0.1$ is $270~{}\mu\rm s$. Comparison with the time for a long-range Rydberg gate shown in Fig. 6 depends on assumptions about the beam pointing time. With sub $\mu\rm s$ beam pointing the motion enabled gate is slower by a factor of about 2 - 5.
IV.3 Lattice surgery
@@
A third approach to implementing gates on pairs of logical qubits is lattice surgery[155]. In order to perform a $\sf CNOT$ gate between two surface code qubits of distance $d$ an auxiliary patch of surface code is utilized. A sequence of $d$ physical gates and $d$ measurements is performed along the boundary of the first logical qubit and the auxiliary region. This merging operation creates a code patch of dimension $d\times 2d$. The merged patch is then split with a sequence of $d$ gates and $d$ measurements between it and the second logical qubit. The net result after $2d$ physical gates and $2d$ measurements is a $\sf CNOT$ gate between the two logical qubits.
@@
A feature of lattice surgery is that all of the operations occur along the boundaries of the logical patches which eliminates the need for long-range gates or long distance motion. The time for the operation if it is performed in place using individual addressing beams that are scanned is $t_{\rm ls}=2d(t_{\sf CZ}+t_{\rm beam}+t_{\rm meas})$, with $t_{\rm meas}$ the measurement time (the time for the additional Hadamards to convert $\sf CZ$ to $\sf CNOT$ is negligible). If the operation is performed by moving the atoms so that there are $d$ closely spaced pairs suitable for a global $\sf CZ$ gate the time is $t_{\rm ls}=2t_{\sf CZ}+4t_{\rm move}+2dt_{\rm meas}.$ The required move distance is a single lattice period $r$. Using the value from the previous section and setting $\delta n=0.1$ for the sequence ($\delta n=0.1/4$ per move) we find $t_{\rm move}=71~{}\mu\rm s$. For a code distance of $d=10$ or more the lattice surgery time is dominated by the measurement time. Referring to Fig. 3 the fastest measurement result that approaches the surface code threshold is data point k in the figure[71] which has $t_{\rm meas}=1~{}\rm ms$. The highest fidelity measurements, which are well below threshold, have $t_{\rm meas}\geq 100~{}\rm ms.$ Even with new decoding strategies[187] that reduce the number of rounds of syndrome measurements from $2d$ to $O(1)$, lattice surgery is not competitive with either long-range Rydberg or transport based gates. Substantial improvements in measurement time while maintaining high fidelity would be needed to change this conclusion.
IV.4 Connectivity comparisons
@@
The previous three sections compared the time needed for $\sf CZ$ gates between surface code logical qubits with code distance $d=10$ using three different approaches. The fastest option is long-range Rydberg interactions, although this does not scale beyond $d=10$ without enhancement of the Rydberg interaction range, which is possible, but challenging. Even at $d=10$ an in-place logical gate is challenging due to the requirement on very small atom spacing within each logical qubit. Achieving high fidelity under these conditions will require high performance optics, capable of producing very small focal spots with low light scattering. For smaller distance codes, say $d=5$, the atom spacing could be increased to a more relaxed $r\sim 2~{}\mu\rm m$ but with our assumed gate fidelity of ${\mathcal{F}}=0.9992$ the achievable circuit depth would be limited to $18001/p_{\rm L}=$ 1800, far short of the Megaquop goal.
@@
The lattice surgery approach is relatively straightforward with current technology, either based on in place Rydberg gates or atom transport. The drawback is the slow speed due to the measurement time bottleneck. This leaves the transport based gates described in Sec. IV.2 as the fastest approach based on current experimental capabilities. The timings indicated in Fig. 8 are for a gate between neighboring surface code patches. Assuming a move distance of $64~{}\mu\rm m$ and a heating increment of $\delta n=0.1$ motional quanta we found a logical gate time of $t_{{\rm L},\sf CZ}=270~{}\mu\rm s$.
@@
$\left\langle t_{{\rm L},\sf CZ}\right\rangle=1.82\times 270=490~{}\mu\rm s.$
It is interesting to compare this with the geometry of Ref. [45] which moved qubit pairs to be entangled to a common zone located on the edge of the qubit array. This has the benefit that the Rydberg beams point to a fixed location and do not have to be scanned. The average Manhattan distance to the $1/3$ power to move to a location just outside the middle of one edge of the array is 1.96, giving $\left\langle t_{{\rm L},\sf CZ}\right\rangle=530~{}\mu\rm s$ which is only slightly larger. In either case we see that the cycle time for the $d=10$, 100 logical qubit array will be at best about a kHz.
V Syndrome extraction and mid-circuit measurements
@@
Two problems arise with this type of measurement. First, although the transition out of the $f_{+}$ level is cycling, there is a non-zero Raman rate for scattering from $f_{+}\rightarrow f_{-}$ via the $|np_{3/2,}f=f_{+}\rangle$ state (e.g. $|6s_{1/2},f=4\rangle\rightarrow|6p_{3/2},f=4\rangle\rightarrow|6s_{1/2},f=3\rangle$ in Cs). To prevent this happening it is possible to measure the state using two $\sigma_{+}$ polarized beams that drive the transition $|ns_{1/2},f_{+},m=f_{+}\rangle\rightarrow|np_{3/2},f_{+}+1,m=f_{+}+1\rangle$ which is cycling and does not suffer from Raman transitions. This works well but leads to transverse heating of the atom since the readout and cooling beams are only applied along a single axis. To mitigate heating loss it is necessary to use relatively deep trap potentials which requires more trap laser power. Details concerning relevant experimental parameters can be found in several papers[63, 62, 111, 64, 65, 188, 68].
@@
The second issue is that for quantum error correction based on qubit measurements for syndrome extraction it is necessary to measure ancilla qubits without disturbing the state of proximal data qubits. The standard measurement approach described above does not fare well because the data qubits act as resonant absorbers with a cross section $\sim\lambda^{2}$ but are only spaced by a few $\lambda$. In order to avoid measurement crosstalk several approaches have been developed based on either transport to a distant measurement zone[77, 45], shelving of data qubits in dark states or hiding of data qubits[189, 145, 190, 70], hiding combined with cavity assisted fast readout[191], or arrays with two atomic isotopes or species[164, 192, 193, 194]. There are also proposals based on electromagnetically induced transparency which have not yet been demonstrated[195, 196] .
@@
From the perspective of speed and fidelity of operations almost all of these techniques require some extra operations, beyond what is required for a state selective measurement without regard to crosstalk. The extra operations are either transport to a cavity or a measurement zone which adds time, or the need to shelve or hide atoms that are not intended to be measured by applying additional control pulses, which adds errors from imperfect operations. The one exception is the two-species approach. When seen as part of a syndrome extraction circuit for error correction the operations required for two-species are the same as would be required when using a single species for data and ancilla qubits. The only significant difference is that the entangling gate is applied between atoms of different types. There are Förster resonances available which make, for example, the interspecies Rb-Cs interaction very strong[164, 197]. Similar resonances can be found in other atomic combinations such as Rb-K[198, 199]. For these reasons a two-species architecture appears very promising although the demonstrated Rydberg entanglement fidelity[112, 200] has not yet reached the same level as for single species gates.
@@
An additional feature of the two-species approach is that the ancillae, which are repeatedly measured, can be recooled without disturbing quantum information stored in the data qubits[193]. Furthermore by swapping the data states to ancilla qubits, the data atoms can also be recooled inside a long calculation which can then continue with the roles of data and ancilla atoms swapped, or the information can be swapped back. This mid-circuit cooling capability appears to be uniquely straightforward in a two-species architecture.
[…]
@@
It is interesting to note that the two-species architecture also provides a path to very fast state measurements without requiring optical cavities. The idea analyzed in [201] is to use the second species as a repetition code to provide $N$ times the photon flux for a code of $N$ atoms. This implies close to $N$ times reduction in the integration time needed to make a measurement. Naively we would like to perform the mapping
@@
where the state $|\psi\rangle$ is copied onto $N$ qubits. This is of course not possible because of the no-cloning theorem[15, 16]. However, we can use a repetition code. If $|\psi\rangle=c_{0}|0\rangle+c_{1}|1\rangle$ then an allowed mapping is
@@
This mapping can be achieved by invoking asymmetric Rydberg interactions[202]. The ancilla qubit to be measured is encoded in one species, say Cs, and the auxiliary $N$ atoms for the repetition code are say Rb atoms, as shown in Fig. 9. Choosing Rydberg states that provide a strong interspecies interaction and a weak intraspecies interaction the mapping can be obtained by exciting state $|1\rangle$ of the ancilla to a Rydberg state that blocks the state transfer of the auxiliary atoms via interspecies blockade. Calculations show that with realistic experimental parameters preparation of the repetition code state can be done in under $5~{}\mu\rm s$ for $N\sim 5$ atoms, which should not be taken as a limit on the possible value of $N$. A related idea using electromagnetically induced transparency with a singe species was demonstrated in [203].
Neutral atom mid-circuit measurements have only been demonstrated relatively recently and are still being actively developed. As such it is too early to say which method will work best at scale, and which atomic element will provide the best performance in terms of highest measurement fidelity, lowest loss, and shortest measurement time. Measurement times for arrays imaged in free space are still orders of magnitude longer than gate times, as can be seen in Fig. 3. Fast cavity assisted measurements have been demonstrated in a few $\mu\rm s$[75, 78], but integrating large atom arrays with cavities remains an outstanding challenge.
VI Outlook

### v3

@@ 4.2 Transport based gates
An alternative approach which enables arbitrary longer range connectivity is qubit motion during circuit operation. The concept of relying on physical motion of qubits dates back to early ideas for scaling trapped ion quantum processors Cirac2000 ; Kielpinski02 and has been implemented with great success in recent trapped ion demonstrations Pino2021 ; Moses2023 . In trapped ion platforms, and also extensions to shuttling of quantum dots in semiconductor structures Flentje2017 ; Zwerver2023 , motion relies on applying a sequence of voltages to closely spaced electrodes to provide a time dependent force on the charged particle. A drawback of the trapped ion approach, in addition to the complexity of controlling many electrodes, is that motion induces substantial heating of the center of mass motional state so recooling is necessary as part of the control sequence.
@@
Neutral atoms trapped in optical potentials defined by light can be transported much more efficiently in terms of the number of required control signals: either by varying the phase of a single optical beam to continuously move the nodes of an interference pattern Mandel2003b , or by scanning an optical tweezer using a varying frequency applied to an acousto-optical deflector Beugnon2007 . When applied to moving arrays of atoms in parallel the optical tweezer approach becomes a powerful technique for connecting logical qubits in a neutral atom processor Bluvstein2022 ; Bluvstein2024 ; Reichardt2024 .
@@
We may compare the time to implement a logical CZ gate using atom motion with the case of long-range Rydberg interactions discussed in Sec. 4.1. The attainable transport speed is limited by the need to minimize motional heating Hickman2020 ; Hwang2025 . Several different transport profiles have been demonstrated including constant jerk Bluvstein2022 , and minimal jerk LLiu2019 ; Matthies2024 ; Finkelstein2024 , where jerk is the time derivative of acceleration Hogan1984 . Remarkably it is also possible to throw and catch atoms over distances of several $\mu\rm m$ Hwang2023 .
@@
For the minimal jerk profile we find using the analysis in Carruthers1965 that the time to move a distance $R$ while incurring $\delta n$ quanta of motional excitation is Chinnarasu2025
@@
Figure 8: Sequence for a transport based logical $\sf CZ$ involves (top to bottom) moving the data qubits a distance $R$, applying a Rydberg pulse, and then moving the data qubits back. The contour plot shows the minimal jerk time to implement the move back and forth, neglecting any additional time for the Rydberg pulse, as a function of the distance $R=dr$, and the round trip motional heating incurred. The inset shows the one-dimensional atomic temperature as a function of the mean number of motional quanta assuming a trap frequency of $\omega_{0}=2\pi\times 100\penalty 10000\ \rm kHz$ and the expression $k_{\rm B}T=\hbar\omega_{0}/\ln(1+1/\langle n\rangle)$.
@@
Parallel implementation of a transport based transversal gate relies on positioning the atom pairs at a small spacing, relative to the distance to neighboring pairs. Doing so allows multiple gate pairs to be performed in the same time step. As long as the array pitch $r$, which is the distance between pairs, satisfies $r\gg r_{g}$, with $r_{g}$ the distance between atoms in one gate pair there will be minimal crosstalk. The sensitivity of the time-optimal $\sf CZ$ gate to interaction crosstalk was studied in Poole2025a where it was found that provided $\eta_{i}=\sum_{j}\Delta_{ij}t_{\rm gate}/(2\pi)<0.01$ the gate infidelity due to crosstalk is upper bounded by $10^{-4}$. Here $\Delta_{ij}$ is the energy shift of Rydberg atom $i$ due to all other atoms $j$, and $t_{\sf CZ}$ is the gate time. In the van der Waals interaction limit the shifts scale as $\Delta_{ij}\simeq\frac{V(r_{g})}{\hbar}\left(\frac{r_{g}}{r_{ij}}\right)^{6}.$ For a high fidelity time-optimal gate we require Jandura2022 $\Omega t_{\sf CZ}/\hbar\geq 7.6$ with $\Omega$ the Rydberg excitation Rabi frequency. Using $(V/\hbar)/\Omega=20$ which is adequate for a high fidelity gate and approximating $\eta_{i}$ by the worst case of $\eta\simeq 4\Delta(r)t_{\sf CZ}/(2\pi)$ we find
@@
Using $r_{g}=1.4\penalty 10000\ \mu\rm m$ (the same tight array spacing as was assumed in Fig. 6) we require $r=6.44\penalty 10000\ \mu\rm m$ so the transport distance for a $d=10$ code is $R=64\penalty 10000\ \mu\rm m$. Referring to Fig. 8 the total transport time at a heating increment of $\delta n=0.1$ is $270\penalty 10000\ \mu\rm s$. Comparison with the time for a long-range Rydberg gate shown in Fig. 6 depends on assumptions about the beam pointing time. With sub $\mu\rm s$ beam pointing the motion enabled gate is slower by a factor of about 2 - 5.
4.3 Lattice surgery
@@
A third approach to implementing gates on pairs of logical qubits is lattice surgery Horsman2012 . In order to perform a $\sf CNOT$ gate between two surface code qubits of distance $d$ an auxiliary patch of surface code is utilized. A sequence of $d$ physical gates and $d$ measurements is performed along the boundary of the first logical qubit and the auxiliary region. This merging operation creates a code patch of dimension $d\times 2d$. The merged patch is then split with a sequence of $d$ gates and $d$ measurements between it and the second logical qubit. The net result after $2d$ physical gates and $2d$ measurements is a $\sf CNOT$ gate between the two logical qubits.
@@
A feature of lattice surgery is that all of the operations occur along the boundaries of the logical patches which eliminates the need for long-range gates or long distance motion. The time for the operation if it is performed in place using individual addressing beams that are scanned is $t_{\rm ls}=2d(t_{\sf CZ}+t_{\rm beam}+t_{\rm meas})$, with $t_{\rm meas}$ the measurement time (the time for the additional Hadamards to convert $\sf CZ$ to $\sf CNOT$ is negligible). If the operation is performed by moving the atoms so that there are $d$ closely spaced pairs suitable for a global $\sf CZ$ gate the time is $t_{\rm ls}=2t_{\sf CZ}+4t_{\rm move}+2dt_{\rm meas}.$ The required move distance is a single lattice period $r$. Using the value from the previous section and setting $\delta n=0.1$ for the sequence ($\delta n=0.1/4$ per move) we find $t_{\rm move}=71\penalty 10000\ \mu\rm s$. For a code distance of $d=10$ or more the lattice surgery time is dominated by the measurement time. Referring to Fig. 3 the high fidelity measurements which are well below the surface code threshold have $t_{\rm meas}$ of at least a few ms. We see that lattice surgery is not competitive with either long-range Rydberg or transport based gates. Substantial improvements in measurement time while maintaining high fidelity would be needed to change this conclusion.
4.4 Connectivity comparisons
@@
The previous three sections compared the time needed for $\sf CZ$ gates between surface code logical qubits with code distance $d=10$ using three different approaches. The fastest option is long-range Rydberg interactions, although this does not scale beyond $d=10$ without enhancement of the Rydberg interaction range, which is possible, but challenging. Even at $d=10$ an in-place logical gate is challenging due to the requirement on very small atom spacing within each logical qubit. Achieving high fidelity under these conditions will require high performance optics, capable of producing very small focal spots with low light scattering. For smaller distance codes, say $d=5$, the atom spacing could be increased to a more relaxed $r\sim 2\penalty 10000\ \mu\rm m$ but with our assumed gate fidelity of ${\mathcal{F}}=0.9992$ the achievable circuit depth would be limited to $1/p_{\rm L}=1800$, far short of the Megaquop goal.
@@
The lattice surgery approach is relatively straightforward with current technology, either based on in place Rydberg gates or atom transport. The drawback is the slow speed due to the measurement time bottleneck. This leaves the transport based gates described in Sec. 4.2 as the fastest approach based on current experimental capabilities. The timings indicated in Fig. 8 are for a gate between neighboring surface code patches. Assuming a move distance of $64\penalty 10000\ \mu\rm m$ and a heating increment of $\delta n=0.1$ motional quanta we found a logical gate time of $t_{{\rm L},\sf CZ}=270\penalty 10000\ \mu\rm s$.
@@
$\left\langle t_{{\rm L},\sf CZ}\right\rangle=1.82\times 270=490\penalty 10000\ \mu\rm s.$
It is interesting to compare this with the geometry of Ref. Bluvstein2024 which moved qubit pairs to be entangled to a common zone located on the edge of the qubit array. This has the benefit that the Rydberg beams point to a fixed location and do not have to be scanned. The average Manhattan distance to the $1/3$ power to move to a location just outside the middle of one edge of the array is 1.96, giving $\left\langle t_{{\rm L},\sf CZ}\right\rangle=530\penalty 10000\ \mu\rm s$ which is only slightly larger. In either case we see that the cycle time for the $d=10$, 100 logical qubit array will be at best about a kHz.
5 Syndrome extraction and mid-circuit measurements
@@
Two problems arise with this type of measurement. First, although the transition out of the $f_{+}$ level is cycling, there is a non-zero Raman rate for scattering from $f_{+}\rightarrow f_{-}$ via the $|np_{3/2,}f=f_{+}\rangle$ state (e.g. $|6s_{1/2},f=4\rangle\rightarrow|6p_{3/2},f=4\rangle\rightarrow|6s_{1/2},f=3\rangle$ in Cs). To prevent this happening it is possible to measure the state using two $\sigma_{+}$ polarized beams that drive the transition $|ns_{1/2},f_{+},m=f_{+}\rangle\rightarrow|np_{3/2},f_{+}+1,m=f_{+}+1\rangle$ which is cycling and does not suffer from Raman transitions. This works well but leads to transverse heating of the atom since the readout and cooling beams are only applied along a single axis. To mitigate heating loss it is necessary to use relatively deep trap potentials which requires more trap laser power. Details concerning relevant experimental parameters can be found in several papers Gibbons2011 ; Fuhrmanek2011 ; Jau2016 ; Kwon2017 ; Martinez-Dorantes2017 ; Martinez-Dorantes2018 ; Chow2023 .
@@
The second issue is that for quantum error correction based on qubit measurements for syndrome extraction it is necessary to measure ancilla qubits without disturbing the state of proximal data qubits. The standard measurement approach described above does not fare well because the data qubits act as resonant absorbers with a cross section $\sim\lambda^{2}$ but are only spaced by a few $\lambda$. In order to avoid measurement crosstalk several approaches have been developed based on either transport to a distant measurement zone Deist2022 ; Bluvstein2024 , shelving of data qubits in dark states or hiding of data qubits Graham2023b ; Lis2023 ; Norcia2023 ; SMa2023 ; Bluvstein2025 , hiding combined with cavity assisted fast readout BHu2025 , or arrays with two atomic isotopes or species Beterov2015 ; CSheng2022 ; Singh2023 ; CFang2025 . There are also proposals based on electromagnetically induced transparency which have not yet been demonstrated Giraldo2022 ; Saglam2023 .
@@
From the perspective of speed and fidelity of operations almost all of these techniques require some extra operations, beyond what is required for a state selective measurement without regard to crosstalk. The extra operations are either transport to a cavity or a measurement zone which adds time, or the need to shelve or hide atoms that are not intended to be measured by applying additional control pulses, which adds errors from imperfect operations. The one exception is the two-species approach. When seen as part of a syndrome extraction circuit for error correction the operations required for two-species are the same as would be required when using a single species for data and ancilla qubits. The only significant difference is that the entangling gate is applied between atoms of different types. There are Förster resonances available which make, for example, the interspecies Rb-Cs interaction very strong Beterov2015 ; Ireland2024 . Similar resonances can be found in other atomic combinations such as Rb-K Samboy2017 ; Otto2020 . For these reasons a two-species architecture appears very promising although the demonstrated Rydberg entanglement fidelity YZeng2017 ; Anand2024 has not yet reached the same level as for single species gates.
@@
An additional feature of the two-species approach is that the ancillae, which are repeatedly measured, can be recooled without disturbing quantum information stored in the data qubits Singh2023 . Furthermore, by mapping the data qubit states onto ancilla qubits, the data atoms can also be recooled inside a long calculation, which can then continue with the roles of data and ancilla atoms swapped, or the information can be swapped back. This mid-circuit cooling capability appears to be uniquely straightforward in a two-species architecture.
Figure 9: Fast state measurements using a two-species repetition code. (left) Surface code layout with Cs data and ancilla qubits and Rb measurement atoms. (right) Interaction potentials for the Rydberg pair states Rb46s46s–Cs48s48s. Colored points are the relative energies of eigenstates of the diagonalized Hamiltonian (shading is proportional to the eigenstate overlap with the target state). Solid lines are fits used to extract $C_{3}$ and $C_{6}$ coefficients. Inset: Ratio of interspecies Rb-Cs interaction energy to the intraspecies Rb-Rb interaction energy. Adapted from Petrosyan2024 .
@@
It is interesting to note that the two-species architecture also provides a path to very fast state measurements without requiring optical cavities. The idea analyzed in Petrosyan2024 is to use the second species as a repetition code to provide $N$ times the photon flux for a code of $N$ atoms. This implies close to $N$ times reduction in the integration time needed to make a measurement. Naively we would like to perform the mapping
@@
where the state $|\psi\rangle$ is copied onto $N$ qubits. This is of course not possible because of the no-cloning theorem Wootters1982 ; Dieks1982 . However, we can use a repetition code. If $|\psi\rangle=c_{0}|0\rangle+c_{1}|1\rangle$ then an allowed mapping is
@@
This mapping can be achieved by invoking asymmetric Rydberg interactions Saffman2009b . The ancilla qubit to be measured is encoded in one species, say Cs, and the auxiliary $N$ atoms for the repetition code are say Rb atoms, as shown in Fig. 9. Choosing Rydberg states that provide a strong interspecies interaction and a weak intraspecies interaction the mapping can be obtained by exciting state $|1\rangle$ of the ancilla to a Rydberg state that blocks the state transfer of the auxiliary atoms via interspecies blockade. Calculations show that with realistic experimental parameters preparation of the repetition code state can be done in under $5\penalty 10000\ \mu\rm s$ for $N\sim 5$ atoms, which should not be taken as a limit on the possible value of $N$. A related idea using electromagnetically induced transparency with a singe species was demonstrated in WXu2021 .
Neutral atom mid-circuit measurements have only been demonstrated relatively recently and are still being actively developed. As such it is too early to say which method will work best at scale, and which atomic element will provide the best performance in terms of highest measurement fidelity, lowest loss, and shortest measurement time. Measurement times for arrays imaged in free space are still orders of magnitude longer than gate times, as can be seen in Fig. 3. Fast cavity assisted measurements have been demonstrated in a few $\mu\rm s$ Bochmann2010 ; Grinkemeyer2025 , but integrating large atom arrays with cavities remains an outstanding challenge.
6 Outlook